OpenStudy (linn99123):

Which ordered pair is a solution to the equation 4x + 7y = –5? A. (–4, –3) B. (–4, 3) C. (4, –3) D. (4, 3)

OpenStudy (phebe):

i would say B but that might be wrong is it @VeritasVosLiberabit

OpenStudy (anonymous):

B is incorrect

OpenStudy (phebe):

ok yaya im rit

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

pay close attention to signs here you want the equation to be -5=-5

OpenStudy (anonymous):

(x,y) try to just plug them into the x and y values directly and solve by trial and error

OpenStudy (anonymous):

I agree with C

OpenStudy (phebe):

C is wrong its B

OpenStudy (phebe):

take another look at it

OpenStudy (anonymous):

Agree with C also. Do not agree on how the answer was given

OpenStudy (anonymous):

@Phebe B give 5=-5 which is untrue

OpenStudy (anonymous):

C is the correct answer given

OpenStudy (anonymous):

@Phebe you are doing something wrong :/

OpenStudy (phebe):

well ok tha is ritsooo its C i guess explain thoee

OpenStudy (anonymous):

look at the problem closely. they want -5 not 5

OpenStudy (phebe):

yea i do do something wrong @Shay17

OpenStudy (anonymous):

4x + 7y = –5? plug in 4,-5 4(4)+7(-3)=-5 16 - 21 =-5 -5=-5 GOSHHH its C

OpenStudy (phebe):

okkkkkkkkkkk

OpenStudy (linn99123):

hits self with pan*

OpenStudy (anonymous):

Can i do it for you ? :P

OpenStudy (linn99123):

Gives reaper the pan*

OpenStudy (anonymous):

xD get back to work

OpenStudy (linn99123):

D: